OpenStudy (anonymous):

Convert to a common log: log3 4

OpenStudy (xapproachesinfinity):

to change the base all you do is this: \(\huge \log_ab=\frac{\ln b}{\ln a}\)

OpenStudy (xapproachesinfinity):

the common log is the natural logarithm ln

OpenStudy (anonymous):

so the answer is 2log3(2)

OpenStudy (xapproachesinfinity):

didn't you see what i did!

OpenStudy (anonymous):

???